Consider a game of “ladder climbing”. There are 5 levels in the game, level 1 is the lowest (bottom) and level 5 is the highest (top). A player starts at the bottom. Each time, a fair coin is tossed. If it turns up heads, the player moves up one rung. If tails, the payer moves down to the very bottom. Once at the top level, the player moves to the very bottom if a tail turns up, and stray at the top if head turns up.

  1. Find the transition probability matrix.
  2. Find the two-step (p2) transition probability matrix.
  3. Find the steady-state distribution of the Markov chains.
Answer #1

Transitio o probability matrix According To nature of game P = 0.5 0.5 0.5 0.5 0.5 0.5 0 0 0 0 o ol 0.5 o o 0 0.5 o 0 o 0. 51Steady-state distribution of markou-chain 0-0.5 C1+12+13 +16 +15) ETI 0.5 l, = 12 - © 0.5 12 -1 - (3) 0.5 13 - 16 - 0.5 14 +0О. 25 xo.5 - 03 П3 = 0.125 П4 = 0.5xo-) 25 П. – О.06 2.5 Пs - O.5 04 +0. 5 Пs Јls - 0.5 x0.06 25 +0.5 x 5 :. Пs = 0.03 125 +0